# Review of nonflow estimation methods and uncertainties in relativistic heavy-ion collisions

### v2

@@ 2.3 Data-driven fitting methods
The away-side dijet correlations are always the most notorious nonflow to handle. If the away-side correlations in $\Delta\eta$ are just like flow correlations that both are relatively uniform, then 2D fitting cannot distinguish them. One may resort to 2D fitting if the away-side correlations are relatively sharp-peaked in $\Delta\phi$ such that one is reasonably confident that it is nonflow. If it is broadly distributed, then there is really no way to distinguish it from flow as Fourier components can describe any functional shape. However, it is more likely that the away-side correlations in $\Delta\eta$ differ from flow; for instance, away-side jet correlations could be more peaked at $\Delta\eta=0$ at relatively high $p_{T}$, or at low $p_{T}$ perhaps dipped at $\Delta\eta=0$. As long as the away-side correlations in $\Delta\eta$ are different from flow, then it is possible to distinguish it in 1D fitting by examining the data structure in $V_{2}(\Delta\eta)$.
@@ 2.3.3 Symmetry method in $(\eta_{1},\eta_{2})$
STAR performed another data-driven analysis by examining $V_{2}(\eta_{1},\eta_{2})$ as a function of the two particles’ pseudorapidities [68]. This method exploits the $\eta$ reflection symmetry in symmetric heavy-ion collisions by comparing two pairs, one at $(\eta_{1},\eta_{2})$ and the other at $(\eta_{1},-\eta_{2})$, and thus does not assume any particular shapes for flow as functions of $\eta$. A $\Delta\eta$-dependent component and a $\Delta\eta$-independent component are identified in $V_{2}(\eta_{1},\eta_{2})$. The former is associated with the combination of nonflow and $\Delta\eta$-dependent flow fluctuations. Linear fits are performed to data as functions of $2\eta_{2}$ and the linear fit parameters are examined as functions of $\eta_{1}$. This leads to a refined functional form as the sum of an exponential and a Gaussian in $\Delta\eta$ to describe the data well. The $\Delta\eta$-independent part is associated with flow plus $\Delta\eta$-independent flow fluctuations. They are further found to be independent of $\eta$ within the limited STAR acceptance of $|\eta|<1$. STAR has also supplemented the analysis with the four-particle cumulant as a function of the particles’ pseudorapidities, albeit large uncertainties, to separate flow fluctuations from the average flow magnitude.
@@
This method has the least model assumption and the functional dependencies on $\Delta\eta$ and $\eta$ are determined from data. The disadvantages of the method are the required high statistics–because $V_{2}(\eta_{1},\eta_{2})$ and especially the four-particle cumulant $V_{2}$ are analyzed in 2-dimensions–and subsequently the large uncertainties, and the reduction of data to the functional forms to describe the various components of flow and nonflow requires a high level of attention. The estimated nonflow in central Au+Au collisions from this method is on the order of 20% with relatively large uncertainty [68].
@@ 3 Discussions and Summary
Nonflow includes all few-body correlations in a collision event except the global flow correlations where all particles are correlated over the entire event. Those nonflow correlations cannot possibly be fully measured, and various estimation/subtraction methods have been devised, as reviewed here in a single place. They can be broadly divided into three categories: (1) $\Delta\eta$-gap, (2) low-multiplicity subtraction, and (3) data-driven fits. All of them have assumptions, some of which are strong and some are less so; they come with different pros and cons. Because of the various degrees of assumptions, assessments of systematic uncertainties on nonflow are challenging.
@@
    The $\Delta\eta$-gap methods, while easy to implement, are not clean. The interpretation of results of a given $\Delta\eta$-gap analysis is subject to issues like the $\Delta\eta$-decorrelation and $\eta$-dependence of flow. The systematic uncertainties are hard to quantify–comparing to full-event (no $\Delta\eta$ gap) benchmark can, for example, be bitten by the usual track-merging detector artifact. An improvement would be to examine the results as functions of the applied $\Delta\eta$-gap size, which would constitute the data-driven 1D fitting method.
@@
    The low-multiplicity subtraction methods come with the large uncertainty in the assumptions of the multiplicity/centrality dependence of nonflow and the arbitrariness in the choice of low-multiplicity event class. Strictly confining within a given set of assumptions, one may arrive at a nonflow estimate with relatively small systematic uncertainty. However, this is only valid when the assumptions are correct; more likely, the estimated nonflow is systematically biased, and loosing the assumptions would yield a wide range of uncertainties.
@@
    The data-driven fitting methods are probably the best at our disposal with the fewest assumptions and least model dependency. However, it requires careful study of data and identification of nonflow correlation shapes. The fitting is often tedious and requires thorough attention and systematic assessment.
In general, nonflow contamination is severe in peripheral collisions and become less so towards more central collisions. However, since the elliptic flow also decreases with increasing centrality because of the more spherical collision zone, nonflow contamination in central collisions can still be appreciable, whereas it is generally the smallest in midcentral collisions. In the top 5% central Au+Au collisions at RHIC, for example, the data-driven fitting methods indicate a nonflow fraction of 20% with a typical relative systematic uncertainty of 20% [68, 69, 63, 62]. From low-multiplicity subtraction, one can get a wide range of nonflow estimate, probably 10–30%, depending on assumptions of multiplicity dependence of nonflow and what low-multiplicity events are taken as nonflow baseline (see Sect. 2.2.1). The former depends on many factors, including a faster increase in the number of nonflow sources than the multiplicity and a change in the correlation of each nonflow source with multiplicity/centrality.
Nonflow subtraction is particularly challenging in small system collisions to search for signs of collectivity in high-multiplicity events. The reasons are many-fold: the nonflow contamination is high and likely dominates the $V_{2}$ measurement, multiplicity selection biases are significant for both low- and high-multiplicity events, how nonflow varies from low- to high-multiplicity collisions is largely unknown partially because of the selection biases. The task is easier at the LHC than at RHIC because of the larger detector longitudinal acceptances, the larger multiplicities produced, and the more likelihood to have collectivity at the higher energies of LHC. At RHIC the analysis is significantly more difficult; for example, the difference between PHENIX [70, 71] and STAR [18, 19] on flow in p+Au, d+Au, and 3He+Au is not fully settled.

### v4

@@ 2.3 Data-driven fitting methods
The away-side dijet correlations are always the most notorious nonflow to handle. If the away-side correlations in $\Delta\eta$ are just like flow correlations such that both are relatively uniform, then 1D fitting cannot distinguish them. One may resort to 2D fitting if the away-side correlations are relatively sharp-peaked in $\Delta\phi$ such that one is reasonably confident that it is nonflow. If it is broadly distributed, then there is really no way to distinguish it from flow as Fourier components can describe any functional shape. However, it is more likely that the away-side correlations in $\Delta\eta$ differ from flow; for instance, away-side jet correlations could be more peaked at $\Delta\eta=0$ at relatively high $p_{T}$ (at the highest-$p_{T}$ extreme the dijet will have to be both at midrapidity), or at low $p_{T}$ perhaps dipped at $\Delta\eta=0$ (e.g. because of longitudinal flow). On the other hand, if the away-side correlations in $\Delta\eta$ are sufficiently different from flow, then it is possible to distinguish it in 1D fitting by examining the data structure in $V_{2}(\Delta\eta)$.
It is noteworthy that the 1D fitting result and the $\Delta\eta$-gap results are certainly related because they both come from the data. However, the flow components extracted from these two methods differ because of the different assumptions involved. The flow component from the $\Delta\eta$-gap method is whatever is measured in the $V_{2}$ beyond the $\Delta\eta$ gap, while that from the 1D fitting method is more sophisticated, dependent of the flow shape and decorrelation in $\Delta\eta$. Nevertheless, one can calculate/integrate from the 1D fitting result the flow component that would be extracted with the $\Delta\eta$-gap method.
@@ 2.3.3 Symmetry method in $(\eta_{1},\eta_{2})$
STAR performed another data-driven analysis by examining $V_{2}(\eta_{1},\eta_{2})$ as a function of the two particles’ pseudorapidities [125]. This method [129] exploits the $\eta$ reflection symmetry in symmetric heavy-ion collisions by comparing two pairs of pseudorapidity bins, one at $(\eta_{1},\eta_{2})$ and the other at $(\eta_{1},-\eta_{2})$, and thus does not assume any particular shapes for flow as functions of $\eta$. Any difference between the two pairs must arise from $\Delta\eta$-dependent physics. The $\Delta\eta$-dependent component and the $\Delta\eta$-independent component can be identified in $V_{2}(\eta_{1},\eta_{2})$. The former is associated with the combination of nonflow and $\Delta\eta$-dependent flow fluctuations (decorrelation). The $\Delta\eta$-independent part is associated with flow plus $\Delta\eta$-independent flow fluctuations. They are found to be independent of $\eta$ within the limited STAR acceptance of $|\eta|<1$. STAR has also supplemented the analysis with the four-particle cumulant $V_{2}\{4\}(\eta_{1},\eta_{2})$ as a function of the particles’ pseudorapidities, albeit large uncertainties, to separate flow fluctuations from the average flow magnitude.
@@
This method has the least model assumption and the functional dependencies on $\eta$ and $\Delta\eta$ are determined from data. The disadvantages of the method are the required high statistics, [^5] The four-particle cumulant $V_{2}\{4\}(\eta_{1},\eta_{2})$ is also analyzed in two dimensions and is limited by statistics with large uncertainties. and the reduction of data to functional forms to economically describe the various components of flow and nonflow requires iterations of trial and error. The estimated nonflow in the most central 0–5% Au+Au collisions from this method is on the order of 20% with a relatively large uncertainty [125].
3 An example case study
As an example, we illustrate the various nonflow subtraction methods using STAR data of dihadron $(\Delta\eta,\Delta\phi)$ correlation functions in Au+Au  collisions at $\sqrt{s_{\textsc{nn}}}=200$ GeV published in Ref. [105]. The data are presented in the form of $\Delta\rho/\sqrt{\rho_{\rm ref}}=\sqrt{\rho_{\rm ref}}(r-1)$ where $r$ is the $(\Delta\eta,\Delta\phi)$ correlation function divided by the properly normalized mixed-event one, and $\sqrt{\rho_{\rm ref}}=dN/d\eta/2\pi$.
The $\Delta\eta$-gap method. One can calculate $V_{n}$ from the correlation function $r(\Delta\eta,\Delta\phi)$ by Eq. (4) with certain minimum $\Delta\eta$ gaps. The values of $V_{2}$ for the top 0–5% centrality are listed in Table 1 for various $\Delta\eta$ gaps, together with the relative reduction from the inclusive one (i.e. no $\Delta\eta$ gap requirement). This reduction corresponds to the removed nonflow fraction by each $\Delta\eta$ gap.
It is noteworthy that the correlation functions have already been corrected for the triangle $\Delta\eta$ acceptance via the mixed-event technique [105]. The calculated numerical value of $V_{n}$ is dependent of this acceptance correlation because nonflow is primarily composed of short-range correlations, decreasing with increasing $\Delta\eta$. Because flow is approximately $\eta$ independent, the $\Delta\eta$-acceptance is sometimes uncorrected in experimental analysis of two-particle cumulant $V_{2}$, e.g., in early works of Ref. [72, 130]. Without the triangle acceptance correction, the calculated $V_{2}$ values would be larger. For numerical illustration, we “uncorrect” the STAR data by the perfect triangle acceptance, and tabulate the obtained $V_{2}$ values in Table 1 together with the corresponding nonflow reductions. In this case the nonflow effects are larger.
Table 1: An example case study of nonflow estimations by various methods in top 0–5% Au+Au collisions at $\sqrt{s_{\textsc{nn}}}=200$ GeV using midrapidity $(\Delta\eta,\Delta\phi)$ correlation data from STAR [105]. Quoted values are two-particle cumulant $V_{2}$ (×$\times 10^{-4}$) with statistical uncertainties. The midrapidity multiplicity densities are $dN/d\eta=5.2$, 13.9, and 671 for 84–93%, 74–84% and 0–5% centralities, respectively. The data are corrected for $\Delta\eta$ acceptance via mixed events. The last column “uncorrects” the data using the perfect triangle acceptance.
Method Implementation $\Delta\eta$-acc. corrected $\Delta\eta$-acc. uncorrected
$\Delta\eta$-gap $6.12\pm 0.05$ $6.64\pm 0.03$
$|\Delta\eta|>1.04$ $5.30\pm 0.10$ $-13\%$ $5.45\pm 0.07$ $-18\%$
$|\Delta\eta|>1.52$ $5.08\pm 0.18$ $-17\%$ $5.25\pm 0.14$ $-21\%$
Low-multiplicity subtraction $1/N$ scaling 84–93% $5.67\pm 0.06$ $-7\%$ $6.07\pm 0.04$ $-9\%$
74–84% $5.13\pm 0.06$ $-16\%$ $5.42\pm 0.03$ $-18\%$
$1/N$ scaling 84–93% $5.10\pm 0.08$ $-4\%$ $5.25\pm 0.08$ $-4\%$
$|\Delta\eta|>1.04$
dipole scaled 84–93% $5.04\pm 0.11$ $-5\%$ $5.21\pm 0.09$ $-4\%$
$|\Delta\eta|>1.04$
near-side scaled 84–93% $4.93\pm 0.13$ $-19\%$ $5.12\pm 0.13$ $-23\%$
1D fit exp. + const $5.23\pm 0.15$ $-15\%$ $-21\%$
Gaus. + const $5.48\pm 0.08$ $-10\%$ $-17\%$
two-Gaus. + const $5.29\pm 0.12$ $-14\%$ $-20\%$
[…]
Figure 5: (a) Per-hadron normalized $\Delta\phi$ correlations in central 0–5% and peripheral 84–93% Au+Au collisions at $\sqrt{s_{\textsc{nn}}}=200$ GeV. The Fourier coefficients are written on the plots. (b) $\Delta\phi$ correlation function at $|\Delta\eta|<1.04$ minus that at $21.04<|\Delta\eta|<$ 2 after properly normalized on the away side ($\pi/2<\Delta\phi<3\pi/2$). The Gaussian curves are to guide the eye. (c) $V_{2}(\Delta\eta)$ as a function of $\Delta\eta$ of particle pair. Fit functions are superimposed to model nonflow correlations atop of a pedestal, the underlying flow. The $\Delta\phi$ correlation data points are symmetrized about $\Delta\phi=0$ and $\pi$. Error bars are statistical uncertainties. Data from Ref. [105] by the STAR Collaboration.
The $1/N$ scaling method. The per-hadron normalized correlated yields, $\frac{1}{N_{\rm trig}}\frac{dN}{d\Delta\phi}$, can be obtained from projections of $\sqrt{\rho_{\rm ref}}\times r$ [105]. These correlated yield distributions are shown in Fig. 5(a) for the most central 0–5% and most peripheral 84–93% collisions. The first three Fourier coefficients are written on the corresponding plot. Using the simple $1/N$ scaling method of Eq. (7), one can readily obtain the “nonflow-subtracted” flow anisotropy, $V_{n}^{\rm sub}$. The values of elliptic flow $V_{2}^{\rm sub}$ for the top 0–5% centrality are tabulated in Table 1. As discussed in Sect. 2.1.1, the result depends on what centrality is used as the low-multiplicity event class. The $V_{2}^{\rm sub}$ calculated using the 74–84% centrality (instead of the 84–93% one) for low-multiplicity subtraction is also tabulated in Table 1. The corresponding nonflow reduction is larger, as one would expect. The effects of the $\Delta\eta$ triangle acceptance are also shown in Table 1.
In this low-multiplicity subtraction exercise we have used pair correlations without $\Delta\eta$ gap. One can certainty use the correlations with a minimum $\Delta\eta$ requirement and then further apply low-multiplicity subtraction, as we show in Table 1 as well. The relative nonflow reduction is smaller because the base $V_{2}$ with the $\Delta\eta$ gap has a significant portion of nonflow already removed. However, the nonflow reduction is finite, indicating remaining nonflow beyond the applied $\Delta\eta$ gap. Note that the $\Delta\eta$-acceptance effect is in this case negligible because the nonflow contribution at large $\Delta\eta$ is relatively small and approximately $\Delta\eta$ independent.
We may repeat the exercise using the directly measured $v_{2}$ values from STAR [130] (the triangle acceptance is not corrected in these data). The measured $v_{2}$ in peripheral 70–80% and top 0–5% central Au+Au collisions are $v_{2}\{2\}^{\rm peri}\approx 6.9\%$ and $v_{2}\{2\}^{\rm cent}\approx 2.4\%$, respectively [130]. Assuming the former is all nonflow, the $1/N$ scaling of nonflow would result in a nonflow of $(v_{2}^{\rm peri})^{2}\frac{N_{\rm ch}^{\rm peri}}{N_{\rm ch}^{\rm cent}}\approx 1.7\times 10^{-4}$, i.e. 30% nonflow in $(v_{2}^{\rm cent})^{2}$ of the latter. Choosing 90-100% centrality as the low-multiplicity baseline would yield a significantly smaller nonflow estimate [131]. As discussed in Sect. 2.2, it is possible that the 90-100% (or 84-93%) centrality, which is defined by the event multiplicity, is biased towards too soft interactions, thus underestimating the nonflow to be used as subtraction for central collisions. On the other hand, the higher 70-80% centrality data likely contain flow, and thus the 30% estimate is likely an overestimate of nonflow. It is subject to the uncertainty what peripheral range best represents the underlying nonflow magnitude of those central collision events.
To avoid the peripheral multiplicity bias, minimum bias (MB) $pp$ collisions may be a better reference for nonflow subtraction. STAR has measured accumulative correlation as a function of $p_{T}$, $\langle{\sum_{i}\cos 2(\phi_{p_{T}}-\phi_{i})}\rangle$, where $\phi_{p_{T}}$ is the azimuthal angle of the particle of interest from a given $p_{T}$ bin and $\phi_{i}$ is that of the reference particle. The flow component in the accumulative correlations is $Nv_{2}(p_{T})\bar{v}_{2}$, and the nonflow contribution would be constant under the $1/N$ scaling assumption so the ratio $pp$/Au+Au of the accumulative correlations can be directly related to the nonflow fraction in Au+Au collisions. The accumulative correlations are presented as functions of $p_{T}$ in $pp$ and Au+Au collisions (Fig. 1 in Ref. [132]). The peripheral 80–100% Au+Au collisions are comparable to $pp$, suggesting that these peripheral collisions are dominated by nonflow. In the top 0–5% central collisions, the measured ratio is on average 15%, or 12% after correcting for the occupancy-dependent detector efficiency of 20% from $pp$ to central Au+Au [133]. One may consider that the mean $\langle{p_{T}}\rangle$ in central Au+Au collisions is larger than in $pp$, and because the accumulative correlation is an increasing function of $p_{T}$, the $N$-weighted $pp$/Au+Au ratio (or the ratio of the accumulative correlations at the corresponding $\langle{p_{T}}\rangle$ values) is smaller, approximately 8% [131]. On the other hand, one may argue that the $pp$/Au+Au ratio should be taken at higher $p_{T}$ for $pp$ than for Au+Au because of jet energy loss in the latter, which would yield a larger ratio. These considerations suggest that nonflow estimations by the $1/N$ scaling, even without multiplicity biases, still have large uncertainties because of strong assumptions made in this method.
The dipole scaling method. With the Fourier coefficients from peripheral and central collisions, we can readily apply the dipole scaling method of Eq. (14). It is generally expected that the negative dipole is approximately inversely proportional to $N$, so the $1/N$ scaling and the dipole scaling would give similar results of $V_{n}^{\rm sub}$. Obviously, the correlations shown in Fig. 5(a) without $\Delta\eta$ gap is not suitable for the dipole scaling method because the dipole $V_{1}$ in central collisions is not even negative because of the large near-side nonflow peak at $\Delta\phi\sim 0$. Instead, we apply the dipole scaling method to the correlations with a minimum $\Delta\eta$ gap of 1.04 (the bin edge of a $\Delta\eta$ bin). The $V_{1}$ values for $\Delta\phi$ correlations in $21.04<|\Delta\eta|<$ 2 are $(-1.40\pm 0.07)\times 10^{-2}$ and $(-1.30\pm 0.10)\times 10^{-4}$ for 84–93% and 0–5% centralities, respectively. The $V_{1}$ ratio of $0.0093\pm 0.0009$ compares reasonably well to the inverse multiplicity ratio of $0.0077$. The $V_{2}^{\rm sub}$ by Eq. (14) is shown in Table 1. Again the nonflow reduction is small because the base $V_{2}$ with the $\Delta\eta$ gap is already void of significant nonflow.
The template fit method is not shown in Table 1. It is a close variation from the dipole scaling method as discussed in Sect. 2.2.3 and Sect. 2.2.4. The difference between Eqs. 13 and 14 is the $V_{1}$ ratio in the denominator, which is only on the order of 1%.
The near-side scaling method. Figure 5(b) shows the per-hadron normalized correlated yield difference in $\Delta\phi$ between small- and large-$\Delta\eta$ ranges, $|\Delta\eta|<1.04$ and $21.04<|\Delta\eta|<$ 2, respectively. A slight normalization adjustment is applied before taking the difference, a couple of percent for peripheral collisions and negligible for central collisions, such that the away-side ($\pi/2<\Delta\phi<3\pi/2$) correlated yield is averaged to zero. This correlation difference represents the near-side short-range nonflow correlations under the assumption of $\Delta\eta$-independent collective flows. The ratio of the near-side correlated yield in central collisions over that in peripheral collisions is $Y_{\rm high}/Y_{\rm low}=2.65\pm 0.21$, [^6] The $\Delta\phi=0$ data point in central collisions may be affected by track splitting so the ratio is taken excluding the $\Delta\phi=0$ data points. and this ratio is applied in the near-side subtraction method of Eq. (8) to the scaling of nonflow in addition to the simple multiplicity dilution. The near-side scaled nonflow subtracted $V_{2}^{\rm sub}$ value is shown in Table 1. The nonflow effect is significantly larger than the simple $1/N$ scaling because the near-side correlated yield is significantly larger in central collisions.
The 1D fitting method. The $V_{2}$ as a function of $|\Delta\eta|$ is shown in Fig. 5(c). The $V_{2}$ value with a certain minimum $\Delta\eta$ gap requirement is simply the pair density-weighted integral of this plot over the range beyond the given $|\Delta\eta|$ gap size. Take the $\Delta\eta$-gap idea a step forward would be the 1D fitting method. For good accuracy, one would want to take into consideration other flow measurements that are less vulnerable to nonflow, such as those from ZDC measurements or multi-particle cumulants as discussed in Sect. 2.3. For the illustration purpose here, we assume simply that the flow is independent of $\Delta\eta$ and use various functional forms for nonflow correlations. Specifically, we use exponential, single Gaussian, and double Gaussian functions. The fits are superimposed in Fig. 5(c). The fit pedestals, corresponding to the “real” flow, are written on the plot, and tabulated in Table 1 with the corresponding nonflow reduction. The flow for the $\Delta\eta$-acceptance uncorrected data would be the same as the fit pedestal. The corresponding nonflow reduction is larger as shown in Table 1, because the inclusive $V_{2}$ measurement is larger containing stronger nonflow contribution.
It is clear from our example case study that the various nonflow subtraction methods result in a wide range of nonflow estimations. The nonflow contamination in the top 0–5% central Au+Au collisions from this illustrative exercise is approximately 20%. The absolute range in the nonflow estimation and thus the relative uncertainty on the flow result is on the order of 10%. The nonflow effect is expected to be smaller at the LHC than at RHIC because of the larger multiplicity, thus larger dilution of nonflow. A significant fraction of nonflow may be removed by a large $\Delta\eta$ gap. This is in line with findings at the LHC [134, 135] and RHIC [136]. However, nonflow is not completely absent at $|\Delta\eta|>1$. Low-multiplicity subtractions and fitting methods indicate an additional a few percent (absolute) nonflow contamination beyond $|\Delta\eta|>1$.
@@ 4 Discussions and Summary
Nonflow includes all few-body correlations in a collision event except the global flow correlations where all particles are correlated over the entire event. Those nonflow correlations cannot possibly be fully measured, and various estimation/subtraction methods have been devised, as reviewed here in a single place. They can be broadly divided into three categories: (1) $\Delta\eta$-gap, (2) low-multiplicity subtraction, and (3) data-driven fits. All of them have assumptions, some of which are strong and some are less so; they come with different pros and cons. Because of the various degrees of assumptions, assessments of systematic uncertainties on nonflow are challenging. We summarize in Table 2 the various nonflow estimation methods regarding their assumptions, pros, cons, and the sources of uncertainties involved.
Table 2: Summary of the assumptions, pros, cons, and sources of uncertainties involved in the various nonflow estimation methods.
Method Implementation Assumptions Pros Cons Uncertainty sources
$\Delta\eta$-gap
1. pair $\Delta\eta$ 2. two-subevent 3. multi-subevent
- nonflow is short-ranged
simple
difficult to assess systematic uncertainties
- choice of $\Delta\eta$ gap size (nonflow beyond $\Delta\eta$) - flow dependence on $\eta$ flow decorrelation in $\Delta\eta$
Low-multiplicity subtraction
1. $1/N$ scaling 2. near-side scaling 3. template fit 4. dipole scaling
- low-multiplicity events have no flow - certain multiplicity dependencies for nonflow - template fit differs from dipole scaling only in base normalization
restrained by low-multiplicity correlations
strong model dependencies; near-side correlation analysis can be difficult at low $p_{T}$
- choice of low-multiplicity events - degree of validity of assumptions (nonflow shape can change, near- and away-side may scale differently, nonflow sources may not scale with multiplicity) - multiplicity selection biases, especially in small systems - detector effects (track merging/splitting) is multiplicity dependent
Data-driven fitting
1. 2D fit 2. 1D fit 3. $\eta$-symmetry
- flow and nonflow shapes are different
least model dependent
fitting can be demanding (trial and error)
- functional forms for nonflow correlations - $\Delta\eta$ dependence of flow and flow fluctuations (which can be input from elsewhere)
@@
    The $\Delta\eta$-gap methods, while simple to implement, are not clean. The interpretation of results of a given $\Delta\eta$-gap analysis is subject to issues like the $\eta$-dependence and $\Delta\eta$-decorrelation of flow. The systematic uncertainties are hard to quantify; comparing to full-event (no $\Delta\eta$ gap) benchmark can, for example, be bitten by the usual track-merging/splitting detector artifacts. An improvement would be to examine the results as functions of the applied $\Delta\eta$-gap size, which would constitute the data-driven 1D fitting method.
@@
    The low-multiplicity subtraction methods come with the large uncertainty in the assumptions of the multiplicity/centrality dependence of nonflow and the arbitrariness in the choice of low-multiplicity event class. Strictly confining within a given set of assumptions, one may arrive at a nonflow estimate with relatively small systematic uncertainty. However, this is only valid when the assumptions are correct; more likely, the estimated nonflow is systematically biased, and loosening the assumptions would yield a wide range of uncertainties.
@@
    The data-driven fitting methods are probably the best at our disposal with the fewest assumptions and least model dependency. However, it requires attentive examination of the data and identification of nonflow correlation shapes. The fitting can be demanding and requires thorough inspection.
Naturally, different methods are different in their degrees of efficacy in removing nonflow. The $\Delta\eta$-gap method is used in essentially all modern flow analyses. The low-multiplicity subtraction methods have been actively explored in data analyses, especially in the context of small-system collisions. The fitting methods take more efforts; the methods are not new but are less explored. We have included an example case study of the various nonflow subtraction methods using the top 0–5% central Au+Au data at $\sqrt{s_{\textsc{nn}}}=200$ GeV published by STAR [105].
In general, nonflow contamination is severe in peripheral collisions and become less so towards more central collisions. However, since the elliptic flow also decreases with increasing centrality because of the more spherical collision zone, nonflow contamination in central collisions can still be appreciable, whereas it is generally the smallest in midcentral collisions. In the top 0–5% central Au+Au collisions at RHIC, for example, the data-driven fitting methods indicate a nonflow fraction of 20%, with a typical relative systematic uncertainty of 20% [125, 137, 123, 122]. From low-multiplicity subtraction, one can get a wide range of nonflow estimate, probably 10–30%, depending on assumptions of multiplicity dependence of nonflow and what low-multiplicity events are taken as nonflow baseline. Multiplicity dependence of nonflow involve two factors, one is a possibly faster increase in the number of nonflow sources than the multiplicity, and the other is a possible change in the correlation of each nonflow source with multiplicity/centrality.
Nonflow subtraction is particularly challenging in small system collisions to search for signs of collectivity in high-multiplicity events. The reasons are many-fold: the nonflow contamination is high and likely dominates the $V_{2}$ measurement, multiplicity selection biases are significant for both low- and high-multiplicity events, how nonflow varies from low- to high-multiplicity collisions is largely unknown (partially because of the selection biases). The task is relatively easier at the LHC than at RHIC because of the larger detector longitudinal acceptances, the larger multiplicities produced, and the more likelihood to have collectivity at the higher energies of LHC. At RHIC the analysis is significantly more difficult; for example, the difference between PHENIX [66, 67] and STAR [31, 32] on flow in $p$+Au, $d$+Au, and 3He+Au is not fully settled.
One may naively expect that some of the methods would at least give the lower limit of nonflow (thus upper limit of flow). For example, the $\Delta\eta$-gap method could give a lower limit because one expects residual nonflow to still contribute beyond a certain $\Delta\eta$ gap. This is true if one confines to the physics interpretation of flow and nonflow within the measured $\Delta\eta$ region. If one infers the $\Delta\eta$-gap measurement of flow to, e.g. the entire $|\Delta\eta|>0$ region, then the measured flow at large $\Delta\eta$ can be lower than the true flow over the entire $\Delta\eta$ range because flow and/or flow fluctuation effects can decrease with $\Delta\eta$. This could then be a lower limit of flow and upper limit of nonflow; however, this is really a physics interpretation issue.
@@
We end our topical review of nonflow by a set of recommendations:
  • Use multiple methods (as many as feasible), possibly combine methods, and assess the range of uncertainties in the results;
  • Always strive to use data-driven methods even though they may be more demanding;
  • Employ information and/or constraints from other measurements as much as possible;
  • Use multi-particle cumulants wherever feasible, keeping in mind the different flow fluctuation effects in two- and multi-particle cumulants;
  • For small systems, take extreme care in assessing the possibly severe multiplicity biases in event selections.
We hope our review summarizing the various nonflow estimation methods in a single place is helpful to future researches.